OpenStudy (anonymous):

How do you put the following into an augmented matrix? -4x-4y=16 4x+5y=14

OpenStudy (jdoe0001):

\(\large \begin{array}{rllll} -4x-4y=16\\ 4x+5y=14 \end{array}\implies \begin{bmatrix} -4&-4|16\\ 4&5\quad |14 \end{bmatrix}\)

OpenStudy (jdoe0001):

you pretty much just grab the coefficients, and stick a line before the last one

OpenStudy (anonymous):

Ok! Then how would you solve it?

OpenStudy (loser66):

@jdoe0001

OpenStudy (jdoe0001):

solve it?

OpenStudy (anonymous):

How would I get the answer?

OpenStudy (jdoe0001):

hmmm the values for "x" and "y"? you'd first need to get the inverse matrix for it, and multiply times the last column

OpenStudy (anonymous):

So would I go for the top line -4+(4)(4) -4+(4)(5) 16+(4)(14)?